As an investor in the equity market you become aware of investment opportunity in 3

corporates. Assuming that the cost of equity is 8%, compute the fair value of X Limited,

Y Limited and Z Limited using the Dividend discount model.


a. Given the history of the company, X Limited is expected to pay a uniform dividend of Rs. 5.00 per share.


b. Being in the IT Industry, Y Limited is expected to pay a dividend of Rs. 4.00 per share

and an increase of 5% year on year thereafter.


c. A pharmaceutical company, Z Limited has been paying a dividend of Rs. 2.00 over for

the last 3 years. The company is expected to do extremely well and increase the dividend

pay-out by 7% year on year. This year the dividend expected is Rs. 8.00.


Additionally, what is the biggest lacuna in the Dividend Discount Model in valuing

stocks? Give an example to explain.


1
Expert's answer
2021-09-14T04:48:33-0400

Cost of Equity = 8%

 

Part A

Dividend per share = 5

 

Part B

Dividend per share expected = 4

Constant Growth rate = 5%

 

Part C

Last Dividend = 2

Expected Dividend = 8

Constant Growth rate = 7%

 Solution

Part A

Constant Dividend

Value of stock = Dividend per share ÷ Cost 

of Equity

"=\frac{5}{8\%}\\\frac{5}{0.08}\\

= 62.50"


Part B

Using Constant Growth Model,

Value of stock = Dividend per share

 expected ÷(Cost of Equity − Growth Rate)

"=\\frac{4 }{(0.08\u2212\u20090.05)}\\\\\n\n= 133.33"



Part C

Using Constant Growth Model,

Value of stock = Dividend per share 

expected ÷(Cost of Equity − Growth Rate)

"= \\frac{8 }{ (0.08-\u20090.07)}\\\\\n\n= 800.00"

stocks based on the net present value of the future dividends.